OpenStudy (anonymous):

Is the given point interior, exterior, or on the circle (x + 2)^2 + (y - 3)^2 = 81? R(6, 3)

OpenStudy (anonymous):

and this one Is the given point interior, exterior, or on the circle (x + 2)2 + (y - 3)2 = 81? Q(3, 0)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

you know what the center and radius of \((x+2)^2+(-33)^2=81\) is ?

OpenStudy (misty1212):

oops \[(x+2)^2+(y-3)^2=81\] is what i meant

OpenStudy (anonymous):

no it doesnt say

OpenStudy (misty1212):

lol yeah i know it doesn't the question is, do YOU know it ? form the equation you are told the center and radius

OpenStudy (misty1212):

\[(x-h)^2+(y--k)^2=r^2\] center is \((h,k)\) and radius is (r\)

OpenStudy (misty1212):

you have \(r^2=81\) so what is \(r\) ?